OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair points. (-2,-12) and (-3,-20)

OpenStudy (anonymous):

the formula for slope is \[\frac{y_2 - y_1}{x_2 - x_1}\] are you familiar with that formula?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I AM GOING TO SET THE PROB

OpenStudy (anonymous):

HOW IS IT WRITTEN IN A SIMPLIFIED FRACTION

OpenStudy (anonymous):

and how would I know if it is undefined?

OpenStudy (anonymous):

@Hero it should be 20 - 12 not 20 + 12

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

how did you figured that compared to their workout?

hero (hero):

use y = mx + b -20 = -3m + b -12 = -2m + b b = 3m - 20 b = 2m - 12 b = b 3m - 20 = 2m - 12 3m - 2m = 20 - 12 m = 8

OpenStudy (anonymous):

he just made a simple error 3m - 20 = 2m - 12 if he adds 20 to both sides and subtracts 2m from both sides it becomes 3m - 2m = 20 - 12

hero (hero):

@abb50, I'm trying to cover my own tracks here. I appreciate you pointing out the mistake, but I already corrected it.

OpenStudy (anonymous):

ok thank you and how could i solve if it is undefined or does it has a fraction?

hero (hero):

If it is undefined, then just write "slope is undefined"

OpenStudy (anonymous):

I think it has a fraction

hero (hero):

Nope, an undefined slope cannot be written in fraction form because it is mathematically undefined, meaning no mathematical expression can represent something that cannot be explained mathematically. In other words, an undefined slope does not exist in math.

OpenStudy (anonymous):

oh ok so this one does not have a fraction and it is undefined

OpenStudy (anonymous):

thank you

hero (hero):

This one has a slope m = 8

OpenStudy (anonymous):

ok so how would i type this as a integer should i just put m=8 for the slope?

hero (hero):

You sound like you're completely lost. If that's the case, I cannot help you.

OpenStudy (anonymous):

lol